# Sharp non-uniqueness for the Navier–Stokes equations in scaling critical spaces

### v1

@@ 4.1. Notations and key lemmas
${\left\|V^{({\rm r},2)}_{j}\right\|}_{\dot{B}_{p,q}^{s}}$ $\leqslant C\lambda_{j}^{-2}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q)\sum_{k\in\mathcal{K}}|a_{k}|^{s}{\left\|\nabla^{2}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}}$ (4.62)
$\quad+C\lambda_{j}^{-2}\lambda_{j-1}^{3-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-n)\ell}}{\sqrt{\ell}}\sum_{k\in\mathcal{K}}|a_{k}|^{s-n}{\left\|\nabla^{2n+2}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}$ (4.63)
$\leqslant C\lambda_{j}^{-2}\lambda_{j-1}^{7-2\mu}\sigma_{j}(s,q)\left(1+\lambda_{j-1}^{(4-\mu)n}\right)$ (4.64)
$\leqslant C\lambda_{j}^{-2}\lambda_{j-1}^{7-2\mu}\sigma_{j}(s,q)$ (4.65)
$\leqslant C\lambda_{j}^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}(s,q).$ (4.66)
This completes the proof. ∎
@@
Step 1. Construction and estimates of the principal stationary flow. Let us define
$\widetilde{\sigma}_{j}(s,q):={\left\|\left\{\sigma_{j}(s,q)\right\}_{j^{\prime}}\right\|}_{\ell^{q}(j^{\prime}\geqslant j)},\qquad j\in\mathbb{N}_{0}.$ (4.67)
It follows from Lemmas 4.8 that
$\sum_{j=1}^{\infty}{\left\|V_{j}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}$ $\leqslant C\sum_{j=1}^{\infty}\lambda_{j-1}^{3-2\mu}\sigma_{j}\left(\frac{n}{p}-1,q\right)$ (4.68)
$\leqslant C\lambda^{3-2\mu}\widetilde{\sigma}_{1}\left(\frac{n}{p}-1,q\right)$ (4.69)
$\leqslant C\lambda^{3-2\mu}\frac{\lambda^{(\frac{n}{p}-1)\mu}}{\sqrt{\mu}}=C\lambda^{3-(3-\frac{n}{p})\mu}.$ (4.70)
@@
${\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\leqslant\sum_{j=1}^{\infty}{\left\|V_{j}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\leqslant C\lambda^{3-(3-\frac{n}{p})\mu}.$ (4.72)
@@
${\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\geqslant{\left\|V_{1}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}-\sum_{j=2}^{\infty}{\left\|V_{j}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}.$ (4.73)
Using (4.38), we have
${\left\|V_{1}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}$ $\geqslant C^{-1}\lambda^{3-2\mu}\sigma_{1}\left(\frac{n}{p}-1,q\right)$ (4.74)
$\geqslant C^{-1}\lambda^{3-2\mu}\frac{\lambda^{(\frac{n}{p}-1)\mu}}{\sqrt{\mu}}=C^{-1}\lambda^{3-(3-\frac{n}{p})\mu}.$ (4.75)
Similarly to (4.68), we see that
$\sum_{j=2}^{\infty}{\left\|V_{j}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}$ $\leqslant C\sum_{j=2}^{\infty}\lambda_{j-1}^{3-2\mu}\sigma_{j}\left(\frac{n}{p}-1,q\right)$ (4.76)
$\leqslant C\lambda_{1}^{3-2\mu}\widetilde{\sigma}_{2}\left(\frac{n}{p}-1,q\right)$ (4.77)
$\leqslant C\lambda_{1}^{3-(3-\frac{n}{p})\mu}.$ (4.78)
Hence, we obtain
${\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\geqslant C^{-1}\lambda^{3-(3-\frac{n}{p})\mu}\left\{1-C^{2}\left(\frac{\lambda}{\lambda_{1}}\right)^{3-(3-\frac{n}{p})\mu}\right\}\geqslant C^{-1}\lambda^{3-(3-\frac{n}{p})\mu}.$ (4.79)
@@
Next, we consider the remainder part. By Lemma 4.8, there holds
$\sum_{j=1}^{\infty}{\left\|V_{j}^{({\rm r})}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ $\leqslant C\sum_{j=1}^{\infty}\lambda_{j}^{-1}\lambda_{j-1}^{3-2\mu}\sigma_{j}\left(\frac{n}{r}-1,1\right)$ (4.80)
$\leqslant C\sum_{j=1}^{\infty}\lambda_{j}^{\frac{n}{r}-2}\lambda_{j-1}^{3-2\mu}$ (4.81)
$\leqslant C\lambda_{1}^{\frac{n}{r}-2}\lambda^{3-2\mu}.$ (4.82)
Thus, due to $r>n/2$, we see that the series
@@
absolutely converges in the strong topology of $\dot{B}_{r,1}^{n/r-1}(\mathbb{R}^{n})$ and
${\left\|V^{({\rm r})}\right\|}_{\dot{B}_{r,1}^{n/r-1}}$ $\leqslant\sum_{j=1}^{\infty}{\left\|V_{j}^{({\rm r})}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}$ (4.84)
$\leqslant C\lambda_{1}^{\frac{n}{r}-2}\lambda^{3-2\mu}=C\lambda^{(1-\frac{n}{p})\mu}\lambda_{1}^{\frac{n}{r}-2}\lambda^{3-(3-\frac{n}{p})\mu}.$ (4.85)
@@
$\begin{aligned} {\left\|V\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}&\leqslant{\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+C{\left\|V^{({\rm r})}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\\ &\leqslant C\left(1+\lambda^{(1-\frac{n}{p})\mu}\lambda_{1}^{\frac{n}{r}-2}\right)\lambda^{3-(3-\frac{n}{p})\mu}\\ &\leqslant C\lambda^{3-(3-\frac{n}{p})\mu},\end{aligned}$ (4.86)
$\begin{aligned} {\left\|V\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}&\geqslant{\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}-C{\left\|V^{({\rm r})}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-1}}\\ &\geqslant C^{-1}\left(1-C\lambda^{(1-\frac{n}{p})\mu}\lambda_{1}^{\frac{n}{r}-2}\right)\lambda^{3-(3-\frac{n}{p})\mu}\\ &\geqslant C^{-1}\lambda^{3-(3-\frac{n}{p})\mu},\end{aligned}$ (4.87)
which implies the optimal estimate (4.1).
@@
we focus on its precise structure. To this end, we may decompose $F$ as
@@
$\lambda_{j-1}^{6}\psi_{j}(x)^{2}\sum_{k\in\mathcal{K}}\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{-6}\mathcal{D}V_{j-1}\right)^{2}k\otimes k$ $=\lambda_{j-1}^{6}\psi_{j}(x)^{2}\left({\rm Id}-\lambda_{j-1}^{-6}\mathcal{D}V_{j-1}\right)$ (4.93)
$=\lambda_{j-1}^{6}\psi_{j}^{2}(x){\rm Id}-\mathcal{D}V_{j-1},$ (4.94)
@@
$V^{({\rm p})}_{j}\otimes V^{({\rm p})}_{j}$={} $2\lambda_{j-1}^{6}\sum_{k_{1},k_{2}\in\mathcal{K}}\Gamma_{j,k_{1}}[V_{j-1}]\Gamma_{j,k_{2}}[V_{j-1}]\Psi_{j,k_{1}}\Psi_{j,k_{2}}k_{1}\otimes k_{2}$ (4.95)
$={}$ $\lambda_{j-1}^{6}\psi_{j}^{2}\sum_{k\in\mathcal{K}}\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{6}\mathcal{D}V_{j-1}\right)^{2}k\otimes k$ (4.96)
$+\frac{\lambda_{j-1}^{6}}{h_{j}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\ell}\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{-6}\mathcal{D}V_{j-1}\right)^{2}\psi_{j}^{2}\cos\left(2\lambda^{\ell}a_{k}\cdot x\right)k\otimes k$ (4.97)
$+2\lambda_{j-1}^{6}\sum_{k_{1}\neq k_{2}}\Gamma_{j,k_{1}}[V_{j-1}]\Gamma_{j,k_{2}}[V_{j-1}]\Psi_{j,k_{1}}\Psi_{j,k_{2}}k_{1}\otimes k_{2}$ (4.98)
$={}$ $\lambda_{j-1}^{6}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1}+F_{1,1;j}+F_{1,2;j},$ (4.99)
@@
$F_{1,1;j}:=\frac{\lambda_{j-1}^{6}}{h_{j}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\ell}\Gamma_{j,k}\left[V_{j-1}\right]^{2}\psi_{j}^{2}\cos\left(2\lambda^{\ell}a_{k}\cdot x\right)k\otimes k,$ (4.100)
$F_{1,2;j}:=2\lambda_{j-1}^{6}\sum_{\begin{subarray}{c}k_{1},k_{2}\in\mathcal{K}\\ k_{1}\neq k_{2}\end{subarray}}\Gamma_{j,k_{1}}[V_{j-1}]\Gamma_{j,k_{2}}[V_{j-1}]\Psi_{j,k_{1}}\Psi_{j,k_{2}}k_{1}\otimes k_{2}.$ (4.101)
@@
$-\Delta V_{j-1}+\mathbb{P}\operatorname{div}\left(\lambda_{j-1}^{6}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1}\right)=-\Delta V_{j-1}+\mathbb{P}\nabla(\lambda_{j-1}^{6}\psi_{j}^{2})+\Delta V_{j-1}=0$ (4.102)
@@
$F_{1}$ $=\sum_{j=1}^{\infty}\left(-\Delta V_{j-1}+\mathbb{P}\operatorname{div}\left(\lambda_{j-1}^{6}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1}+F_{1,1;j}+F_{1,2;j}\right)\right)$ (4.103)
@@
${\left\|F_{1,1;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}$ $\leqslant C\lambda_{j-1}^{6-4\mu}\sum_{\ell\in\lambda_{j}}\frac{\lambda^{(\frac{n}{r}-2)\ell}}{\ell}\sum_{k\in\mathcal{K}}{\left\|\Gamma_{j,k}[V_{j-1}]\psi_{j}\right\|}_{L^{r}}$ (4.106)
$\quad+C\lambda_{j-1}^{6-4\mu}\sum_{\ell\in\lambda_{j}}\frac{\lambda^{-2\ell}}{\ell}\sum_{k\in\mathcal{K}}{\left\|\nabla^{n}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{r}\cap L^{1}}$ (4.107)
$\leqslant C\lambda_{j-1}^{6-4\mu}\cdot C\lambda_{j-1}^{(\frac{n}{r}-2)\mu}\cdot C\lambda^{\frac{n}{r}}+C\lambda_{j-1}^{6-4\mu}\cdot C\lambda_{j-1}^{-2\mu}\cdot C\lambda^{n}\lambda_{j-1}^{2n}$ (4.108)
$\leqslant C\lambda_{j-1}^{2n+6-4\mu}.$ (4.109)
For the estimate of $F_{1,2;j}$, since it is rewritten as
$F_{1,2;j}(x)=\frac{\lambda_{j-1}^{6}}{h_{j}}\sum_{\pm}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\sum_{\begin{subarray}{c}k_{1}\neq k_{2}\end{subarray}}$ $\frac{1}{\sqrt{\ell_{1}\ell_{2}}}\Gamma_{j,k_{1}}[V_{j-1}](x)\Gamma_{j,k_{2}}[V_{j-1}](x)$ (4.110)
@@
${\left\|F_{1,2;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\leqslant{}$ $C\lambda_{j-1}^{6-2\mu}\sum_{\pm}\sum_{\begin{subarray}{c}\ell_{1}\in\Lambda_{j}\end{subarray}}\sum_{\ell_{2}:\ell_{2}\geqslant\ell_{1}}\sum_{\begin{subarray}{c}k_{1}\neq k_{2}\end{subarray}}$ (4.112)
@@
$\left.\quad+\frac{\lambda^{-2\ell_{2}}}{\sqrt{\ell_{1}\ell_{2}}}|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|^{-2}{\left\|\nabla^{n}(\Gamma_{j,k_{1}}[V_{j-1}]\Gamma_{j,k_{2}}[V_{j-1}]\psi_{j}^{2})\right\|}_{L^{r}}\right)$ (4.114)
$\leqslant{}$ $C\lambda_{j-1}^{6-4\mu}\cdot C\lambda_{j-1}^{(\frac{n}{r}-2)\mu}\cdot C\lambda^{\frac{n}{r}}+C\lambda_{j-1}^{6-4\mu}\cdot C\lambda_{j-1}^{-2\mu}\cdot C\lambda^{n}\lambda_{j-1}^{2n}$ (4.115)
$\leqslant{}$ $C\lambda_{j-1}^{2n+6-4\mu},$ (4.116)
@@
$|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|^{2}={}$ $\left(\lambda^{\ell_{1}-\ell_{2}}|a_{k_{1}}|\pm|a_{k_{2}}|\cos\vartheta(a_{k_{1}},a_{k_{2}})\right)^{2}+|a_{k_{2}}|^{2}\sin^{2}\vartheta(a_{k_{1}},a_{k_{2}})$ (4.117)
$\geqslant{}$ $|a_{k_{2}}|^{2}\sin^{2}\vartheta(a_{k_{1}},a_{k_{2}})>0.$ (4.118)
@@
$\leqslant C\sum_{j=1}^{\infty}\lambda_{j-1}^{2n+6-4\mu}$ (4.120)
$\leqslant{}C\lambda^{2n+6-4\mu}=C\lambda^{-\beta}.$ (4.121)

### v2

@@ 4.1. Notations and key lemmas
${\left\|V^{({\rm r},2)}_{j}\right\|}_{\dot{B}_{p,q}^{s}}$ $\leqslant C\lambda_{j-1}^{4-2\mu}\sigma_{j}(s-1,q)\sum_{k\in\mathcal{K}}{\left\|\nabla(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}}$ (4.64)
$\quad+C\lambda_{j-1}^{4-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-1-n)\ell}}{\sqrt{\ell}}\sum_{k\in\mathcal{K}}{\left\|\nabla^{2n+1}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}$ (4.65)
$\leqslant C\lambda^{nj}\lambda_{j-1}^{6-2\mu}\sigma_{j}(s-1,q)\left(1+\lambda_{j-1}^{(4-\mu)n}\right)$ (4.66)
$\leqslant C\lambda^{nj}\lambda_{j-1}^{6-2\mu}\sigma_{j}(s-1,q)$ (4.67)
and
${\left\|V^{({\rm r},3)}_{j}\right\|}_{\dot{B}_{p,q}^{s}}$ $\leqslant C\lambda_{j-1}^{4-2\mu}\sigma_{j}(s-2,q)\sum_{k\in\mathcal{K}}{\left\|\nabla^{2}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}}$ (4.68)
$\quad+C\lambda_{j-1}^{4-2\mu}\sum_{\ell\in\Lambda_{j}}\frac{\lambda^{(s-2-n)\ell}}{\sqrt{\ell}}\sum_{k\in\mathcal{K}}{\left\|\nabla^{2n+2}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{p}\cap L^{1}}$ (4.69)
$\leqslant C\lambda^{nj}\lambda_{j-1}^{8-2\mu}\sigma_{j}(s-2,q)\left(1+\lambda_{j-1}^{(4-\mu)n}\right)$ (4.70)
$\leqslant C\lambda^{nj}\lambda_{j-1}^{8-3\mu}\sigma_{j}(s-1,q).$ (4.71)
Gathering (4.63), (4.67), and (4.71), we obtain (4.42) and complete the proof. ∎
@@
Step 1. Construction and estimates of the principal stationary flow. It follows from Lemmas 4.9 that
${\left\|V_{j}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}$ $\leqslant C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\sigma_{j}\left(\frac{n}{p}-1,q\right)$ (4.72)
$\leqslant C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-2\mu}\frac{\lambda_{j-1}^{(\frac{n}{p}-1)\mu}}{\sqrt{\mu\log_{\lambda}\lambda_{j-1}}}\leqslant C\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-(3-\frac{n}{p})\mu}.$ (4.73)
@@
${\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\leqslant\sum_{j=1}^{\infty}{\left\|V_{j}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\leqslant C\lambda^{\frac{n}{p}+4-(3-\frac{n}{p})\mu}.$ (4.75)
@@
${\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\geqslant{}$ ${\left\|V_{1}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}-\sum_{j=2}^{\infty}{\left\|V_{j}^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}$ (4.76)
$\geqslant{}$ $C^{-1}\lambda^{\frac{n}{p}}\lambda^{4-2\mu}\sigma_{1}\left(\frac{n}{p}-1,q\right)-C\sum_{j=2}^{\infty}\lambda^{\frac{n}{p}j}\lambda_{j-1}^{4-(3-\frac{n}{p})\mu}$ (4.77)
$\geqslant{}$ $C^{-1}\lambda^{\frac{n}{p}}\lambda^{4-(3-\frac{n}{p})\mu}-C\lambda^{\frac{2n}{p}}\lambda_{1}^{4-(3-\frac{n}{p})\mu}$ (4.78)
$\geqslant{}$ $C^{-1}\lambda^{\frac{n}{p}+4-(3-\frac{n}{p})\mu}$ (4.79)
@@
Next, we consider the remainder part. Let $\rho>n/2$. Then, similarly as above, we see that the series
@@
absolutely converges in the strong topology of $\dot{B}_{\rho,1}^{n/\rho-1}(\mathbb{R}^{n})$ and
${\left\|V^{({\rm r})}\right\|}_{\dot{B}_{\rho,1}^{n/\rho-1}}$ $\leqslant\sum_{j=1}^{\infty}{\left\|V_{j}^{({\rm r})}\right\|}_{\dot{B}_{\rho,1}^{\frac{n}{\rho}-1}}$ (4.81)
$\leqslant C\sum_{j=1}^{\infty}\lambda^{nj}\lambda_{j-1}^{6-2\mu}\sigma_{j}\left(\frac{n}{\rho}-2,1\right)$ (4.82)
$\leqslant C\sum_{j=1}^{\infty}\lambda^{nj}\lambda_{j-1}^{6-(4-\frac{n}{\rho})\mu}$ (4.83)
$\leqslant C\lambda^{n+6-(4-\frac{n}{\rho})\mu}$ (4.84)
$=C\lambda^{n(1-\frac{1}{p})+2-(1+\frac{n}{p}-\frac{n}{\rho})\mu}\lambda^{\frac{n}{p}+4-(3-\frac{n}{p})\mu}.$ (4.85)
@@
${\left\|V\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}$ $\leqslant{\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}+C{\left\|V^{({\rm r})}\right\|}_{\dot{B}_{p,1}^{\frac{n}{p}-1}}$ (4.86)
$\leqslant C\left(1+\lambda^{n(1-\frac{1}{p})+2-\mu}\right)\lambda^{\frac{n}{p}+4-(3-\frac{n}{p})\mu}$ (4.87)
$\leqslant C\lambda^{\frac{n}{p}+4-(3-\frac{n}{p})\mu}=C\lambda^{-\alpha(p)},$ (4.88)
and
${\left\|V\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}$ $\geqslant{\left\|V^{({\rm p})}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}-C{\left\|V^{({\rm r})}\right\|}_{\dot{B}_{p,1}^{\frac{n}{p}-1}}$ (4.89)
$\geqslant C^{-1}\left(1-C^{2}\lambda^{n(1-\frac{1}{p})+2-\mu}\right)\lambda^{\frac{n}{p}+4-(3-\frac{n}{p})\mu}$ (4.90)
$\geqslant C^{-1}\lambda^{\frac{n}{p}+4-(3-\frac{n}{p})\mu}=C^{-1}\lambda^{-\alpha(p)},$ (4.91)
provided that $\lambda$ is chosen sufficiently large. Thus, we obtain (4.1).
@@
we focus on its precise structure. To this end, we decompose $F$ as
@@
$\lambda_{j-1}^{8}\psi_{j}^{2}\sum_{k\in\mathcal{K}}\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{-8}\mathcal{D}V_{j-1}\right)^{2}k\otimes k$ $=\lambda_{j-1}^{8}\psi_{j}^{2}\left({\rm Id}-\lambda_{j-1}^{-8}\mathcal{D}V_{j-1}\right)$ (4.97)
$=\lambda_{j-1}^{8}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1},$ (4.98)
@@
$V^{({\rm p})}_{j}\otimes V^{({\rm p})}_{j}={}2\lambda_{j-1}^{8}\sum_{k_{1},k_{2}\in\mathcal{K}}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\Phi_{j,k_{1},\ell_{1}}\otimes\Phi_{j,k_{2},\ell_{2}}$ (4.99)
$\quad={}\lambda_{j-1}^{8}\psi_{j}^{2}\sum_{k\in\mathcal{K}}\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{-8}\mathcal{D}V_{j-1}(x)\right)^{2}k\otimes k$ (4.100)
$\qquad+\frac{\lambda_{j-1}^{8}}{h_{j}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\ell}\Gamma_{k}\left({\rm Id}-\lambda_{j-1}^{-8}\mathcal{D}V_{j-1}(x)\right)^{2}\psi_{j}^{2}(x)\cos\left(2\lambda^{\ell}a_{k}\cdot x\right)k\otimes k$ (4.101)
$\qquad+2\lambda_{j-1}^{8}\sum_{\begin{subarray}{c}k_{1},k_{2}\in\mathcal{K}\\ k_{1}\neq k_{2}\end{subarray}}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\Phi_{j,k_{1},\ell}(x)\Phi_{j,k_{2},\ell}(x)$ (4.102)
$\quad={}\lambda_{j-1}^{8}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1}+F_{1,1;j}+F_{1,2;j},$ (4.103)
@@
$F_{1,1;j}:=\frac{\lambda_{j-1}^{8}}{h_{j}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\Lambda_{j}}\frac{1}{\ell}\Gamma_{j,k}\left[V_{j-1}\right]^{2}\psi_{j}^{2}\cos\left(2\lambda^{\ell}a_{k}\cdot x\right)k\otimes k,$ (4.104)
$F_{1,2;j}:=2\lambda_{j-1}^{8}\sum_{\begin{subarray}{c}k_{1},k_{2}\in\mathcal{K}\\ k_{1}\neq k_{2}\end{subarray}}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\Phi_{j,k_{1},\ell_{1}}(x)\otimes\Phi_{j,k_{2},\ell_{2}}(x).$ (4.105)
@@
$-\Delta V_{j-1}+\mathbb{P}\operatorname{div}\left(\lambda_{j-1}^{8}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1}\right)=-\Delta V_{j-1}+\mathbb{P}\nabla(\lambda_{j-1}^{8}\psi_{j}^{2})+\Delta V_{j-1}=0$ (4.106)
@@
$F_{1}$ $=\sum_{j=1}^{\infty}\left(-\Delta V_{j-1}+\mathbb{P}\operatorname{div}\left(\lambda_{j-1}^{8}\psi_{j}^{2}{\rm Id}-\mathcal{D}V_{j-1}+F_{1,1;j}+F_{1,2;j}\right)\right)$ (4.107)
@@
${\left\|F_{1,1;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}$ $\leqslant C\lambda_{j-1}^{8-4\mu}\sum_{\ell\in\lambda_{j}}\frac{\lambda^{(\frac{n}{r}-2)\ell}}{\ell}\sum_{k\in\mathcal{K}}{\left\|\Gamma_{j,k}[V_{j-1}]\psi_{j}\right\|}_{L^{r}}$ (4.110)
$\quad+C\lambda_{j-1}^{8-4\mu}\sum_{\ell\in\lambda_{j}}\frac{\lambda^{-2\ell}}{\ell}\sum_{k\in\mathcal{K}}{\left\|\nabla^{n}(\Gamma_{j,k}[V_{j-1}]\psi_{j})\right\|}_{L^{r}\cap L^{1}}$ (4.111)
$\leqslant C\lambda_{j-1}^{8-4\mu}\cdot C\lambda_{j-1}^{(\frac{n}{r}-2)\mu}\cdot C\lambda^{\frac{n}{r}j}+C\lambda_{j-1}^{8-4\mu}\cdot C\lambda_{j-1}^{-2\mu}\cdot C\lambda^{nj}\lambda_{j-1}^{2n}$ (4.112)
$\leqslant C\lambda_{j-1}^{4n+8-4\mu}.$ (4.113)
For the estimate of $F_{1,2;j}$, since this part is rewritten as
$F_{1,2;j}(x)=\frac{\lambda_{j-1}^{8}}{h_{j}}\sum_{\pm}\sum_{\ell_{1},\ell_{2}\in\Lambda_{j}}\sum_{\begin{subarray}{c}k_{1}\neq k_{2}\end{subarray}}$ $\frac{1}{\sqrt{\ell_{1}\ell_{2}}}\Gamma_{j,k_{1}}[V_{j-1}](x)\Gamma_{j,k_{2}}[V_{j-1}](x)$ (4.114)
@@
${\left\|F_{1,2;j}\right\|}_{\dot{B}_{r,1}^{\frac{n}{r}-2}}\leqslant{}$ $C\lambda_{j-1}^{8-4\mu}\sum_{\pm}\sum_{\begin{subarray}{c}\ell_{1}\in\Lambda_{j}\end{subarray}}\sum_{\ell_{2}:\ell_{2}\geqslant\ell_{1}}\sum_{\begin{subarray}{c}k_{1}\neq k_{2}\end{subarray}}$ (4.116)
@@
$\left.\quad+\frac{\lambda^{-2\ell_{2}}}{\sqrt{\ell_{1}\ell_{2}}}|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|^{-2}{\left\|\nabla^{n}(\Gamma_{j,k_{1}}[V_{j-1}]\Gamma_{j,k_{2}}[V_{j-1}]\psi_{j}^{2})\right\|}_{L^{r}\cap L^{1}}\right)$ (4.118)
$\leqslant{}$ $C\lambda_{j-1}^{8-4\mu}\cdot C\lambda_{j-1}^{(\frac{n}{r}-2)\mu}\cdot C\lambda^{\frac{n}{r}}+C\lambda_{j-1}^{8-4\mu}\cdot C\lambda_{j-1}^{-2\mu}\cdot C\lambda^{n}\lambda_{j-1}^{2n}$ (4.119)
$\leqslant{}$ $C\lambda_{j-1}^{4n+8-4\mu},$ (4.120)
@@
$|\lambda^{\ell_{1}-\ell_{2}}a_{k_{1}}\pm a_{k_{2}}|^{2}={}$ $\left(\lambda^{\ell_{1}-\ell_{2}}\pm\cos\vartheta(a_{k_{1}},a_{k_{2}})\right)^{2}+\sin^{2}\vartheta(a_{k_{1}},a_{k_{2}})$ (4.121)
$\geqslant{}$ $\sin^{2}\vartheta(a_{k_{1}},a_{k_{2}})>0.$ (4.122)
@@
$\leqslant C\sum_{j=1}^{\infty}\lambda_{j-1}^{4n+8-4\mu}$ (4.124)
$\leqslant{}C\lambda^{4n+8-4\mu}=C\lambda^{-\beta}.$ (4.125)